OpenStudy (amour.anohs):

Rearrange the formula V = 1/3 πr2h for h.

OpenStudy (greatlife44):

This is very straight forward: \[V = \frac{ 1 }{ 3 }\pi*r^{2}*h \] \[3*V = 3*\frac{ 1 }{ 3 }\pi*r^{2}*h \] \[\frac{ 3*V }{ \pi*r^{2} } = h \]

OpenStudy (amour.anohs):

I Got It, But Thanks @greatlife44